OpenStudy (anonymous):

two consecutive odd negative integers have a product of 483, find the integers

OpenStudy (anonymous):

Let the two integers be x and x+2

OpenStudy (anonymous):

x(x+2)=483 x^2+2x-83=0 x=21 or -23 Taking only negative x=-23 and -23+2=-21